OpenStudy (anonymous):

Find the relative extrema of y=4x+x^-1

OpenStudy (anonymous):

What is extrema ? Maxima or Minima

OpenStudy (anonymous):

maxima

OpenStudy (anonymous):

Ah then Here you go \[\frac{dy}{dx} = 4 - \frac{1}{x^2}\] \[Put\:\:\:\:\:\:\frac{dy}{dx} = 0\] \[4 = \frac{1}{x^2}\] \[x^2 = \frac{1}{4}\] \[x = \pm \

OpenStudy (anonymous):

\[x = \pm \frac {1}{2}\]

OpenStudy (anonymous):

Sorry \[\frac{-1}{2}\]is the maxima

OpenStudy (anonymous):

It is correct WooBin do you wan to ask anything

OpenStudy (anonymous):

umm...i think its better that you put some explanation so i can fully understand your asnwer. but this okay it helps a lot.

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Now when we take the second derivative \[\frac{d^2y}{dx^2} = \frac{2}{x^3}\]

OpenStudy (anonymous):

Then we put the points -1/2 , +1/2

OpenStudy (anonymous):

\[x= \frac{-1}{2}\] \[-16\]------------------->So this is Maxima \[x=+\frac{1}{2}\] \[16\]-------------------->This is Minima

OpenStudy (anonymous):

Now we have -1/2 has maxima point and 1/2 as minima point

OpenStudy (anonymous):

Now woobin do you understand it now ?

OpenStudy (anonymous):

Call me whenever you have some doubts :D

OpenStudy (anonymous):

A visual aid is attached.